OpenStudy (anonymous):

EASY QUESTION Jorge put $500 into a savings account that earns 2% compounded interest annually. So after the first year, he will have $510, after the second year, he will have $520.20, and after the third year, he will have $530.60. If he does not deposit any more money, how many years will it take until Jorge has over $600 in his account? it will take 9 months, right?

OpenStudy (amistre64):

start with the ending balance: 600 = 530.60(1.02)^n

OpenStudy (debbieg):

Are you sure you mean months?

OpenStudy (amistre64):

A = P(r)^n A/P = r^n ln(A/P) = ln(r^n) ln(A/P) = n ln(r) ln(A/P) ------- = n ln(r)

OpenStudy (amistre64):

if we add in the first 3 years ... then 9 years and a few months seems good